given f(x)= -x^3+2x^2-2x-2, find f(-2) and f(0)
Answer by checkley71(8403)

F(X)=-X^3+2X^2-2X-2
F(-2)=-(-2)^3+2(-2)^2-2(-2)+2
F(-2)=-(-8)+2(4)+4+2
F(-2)=8+8+6
F(-2)=22 ANSWER.
F(0)=-(0)^3+2(0)^2-2*0+2
F(0)=0+0-0+2
F(0)=2 ANSWER.